Car rental agency A will rent a compact car for $35 per day and an additional charge of $0.24 per mile. Car rental agency B will charge only $0.16 per mile but charges $41 per day. If Adam wanted to rent a car for three days, how many miles would Adam have to drive to make car rental agency B a better bargain?
Found 2 solutions by checkley77, kim4567:
Answer by checkley77(12569)

3*35+.24x=3*41+.16x
105+.24x=123+.16x
.24x-16x=123-105
.08x=18
x=.18/.08
x=225 miles will make the 2 plans equal.
Proof:
3*35+.24*225=3*41+.16*225
105+54=123+36
159=159

Answer by kim4567(1)